OpenStudy (anonymous):

Solve 4 < 2x - 2 < 10. show all steps Solve 4 < 2x - 2 < 10. show all steps @Mathematics

OpenStudy (anonymous):

add two divide by two

OpenStudy (anonymous):

how

OpenStudy (anonymous):

that doesnt help

OpenStudy (anonymous):

6<2x<12 3<x<6

OpenStudy (anonymous):

4 < 2x - 2 < 10 +2 +2 +2 6 < 2x < 12 3 < x < 6

OpenStudy (anonymous):

\[4<2x-2<10\] \[4+2<2x-2+2<10+2\] \[6<2x<12\] \[\frac{6}{2}<\frac{2x}{2}<\frac{12}{2}\] \[3<x<6\]